A student throws a ball straight down off a very high building with a magnitude of 3 m/s and accelerates in free fall for 2.0 s what is the velocity of the ball at 2.0 s
1
Expert's answer
2020-09-29T09:40:17-0400

The velocity of a ball

"v_f=v_i+at=\rm 3\: m/s+9.8\: m/s^2\times 2\: s=22.6\; m/s"